[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 06, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 6 20:24:14 UTC 2017


HSU Shift Log for Thu Jul 06 2017 16:24:08 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Sometimes when we open menus to load products, the "Accept"  
and "Close" portion of the pop-up window are not visible for us to click  
and load the product. Not sure if this is a known problem and already being  
worked on.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Split shift between Dave R./Mike/Robbie
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170706/32438b0a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list